OpenStudy (anonymous):

I'm puzzled by this operation: a^2 - 2ab + b^2 / a - b; a * a - 2ab + b * b and I get a - 2ab + b which doesn't seem to be right. Can you explain what I may be doing wrong? Thanks!

OpenStudy (dayakar):

\[\left( a-b \right)^{2}\div(a-b)\]

OpenStudy (dayakar):

(a-b)(a-b)/(a-b) =a-b

OpenStudy (dayakar):

is it clear

OpenStudy (anonymous):

Oh, now I understand what I was doing wrong! Thanks! :-)

OpenStudy (dayakar):

ok

OpenStudy (anonymous):

The result is indeed a - b :) which is what I get now, thanks to you! :)